OpenStudy (anonymous):

f(x)=x^(3)+3x, find each value b. f(1/4) c. f(√2)

OpenStudy (anonymous):

(1/4)^3 + (3/4) sqrt(2)^3 + 3*sqrt(2)

OpenStudy (anonymous):

that took alot of work

OpenStudy (anonymous):

\[\left( \sqrt{2} \right)^{3}\]+\[3\sqrt{2}\]right?

OpenStudy (anonymous):

but pls can i see the working?

OpenStudy (anonymous):

for the second one, \[f(\sqrt{2})=(\sqrt{2})^3+3(\sqrt{2})\\= 2^{3/2}+3\times2^{1/2}\]

OpenStudy (anonymous):

pls electrkid how do u work it?i need it in the squaroot form

OpenStudy (anonymous):

i see what \(f(\sqrt{2})\) means is, replace every "x" in f(x) by \(\sqrt{2}\). and that is exactly what you do

OpenStudy (anonymous):

follow?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or better even, \[ f(x)=x^3+3x^2 \] factorize the "x" \[f(x)=x(x^2+3)\\ f(\sqrt{2})=\sqrt{2}[(\sqrt{2})^2+3]\\=\sqrt{2}(2+3)\\=5\sqrt{2}\]

OpenStudy (anonymous):

how do u simplify \[\left( \sqrt{2} \right)^{3}\]?

OpenStudy (anonymous):

I did not have to. I factored the "x"

OpenStudy (anonymous):

but if you want to, then you can do this: \[(\sqrt{2})^3=(\sqrt{2})^2(\sqrt{2})=2\sqrt{2}\]